OpenStudy (anonymous):

Enough of a monoprotic acid is dissolved in water to produce a 0.0121 M solution. The pH of the resulting solution is 6.49. Calculate the Ka for the acid.

OpenStudy (vincent-lyon.fr):

This problem is similar to that I answered yesterday, even less difficult. 1. Write chemical equation of AH with water. 2. pH allows you to know \([H_3O^+]\) and \([A^-]\) 3. deduce \([AH]\) by subtracting from molarity of solution. 4. Use expression of Ka to compute its numerical value.

OpenStudy (cherylim23):

Ka = [H3O+][A-]/[HA] Since pH = 6.49, [H3O+] = 10^(-6.49) [A-] = [H3O+] = 10^(-6.49) and [HA] = 0.0121 Put this into the acidity constant expression and you should get Ka= 8.65 x 10^-12